I'm helping a 4th grader figure out this problem:
The garden has a perimeter of 90 in. And its length is twice as long as width. What is the length and width of the garden?
I can remember the formula and how do I explain it to a 4th grader????
Thank you soooo much for your help!!!!!

Answer by htmentor(1343)

I'm helping a 4th grader figure out this problem:
The garden has a perimeter of 90 in. And its length is twice as long as width. What is the length and width of the garden?
======================================
Let w = the width of the garden
Since the length is twice as long, the length, l = 2w
The perimeter, P = 2l + 2w = 90
90 = 2(2w) + 2w
Solve for w:
4w + 2w = 90
6w = 90
w = 15
So the width is 15 in., the length is 30 in.